OpenStudy (anonymous):

For all Fetha, prove that cos fetha =1/sec fetha

OpenStudy (anonymous):



OpenStudy (anonymous):

umm

OpenStudy (anonymous):

Now, cos theta=b/h........i sec theta=h/b.......ii

OpenStudy (anonymous):

So, b/h=1/sec theta cos theta = 1/ sec theta

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

@fari321 ???

OpenStudy (anonymous):

ummm you have to solve it using the reciprocal identities or quotient identities

OpenStudy (anonymous):

I think I used reciprocal identities

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

So I was wrong

OpenStudy (anonymous):

But I think cos theta =1/sec theta Cant be proved by other ways.

OpenStudy (anonymous):

Let me call the expert @satellite73 can u PLZ help

OpenStudy (anonymous):

look at the image i posted

OpenStudy (anonymous):

I think I did the same

OpenStudy (anonymous):

not sure what you need by definition \(\sec(\theta)=\frac{1}{\cos(\theta)}\) so \(\cos(\theta)=\frac{1}{\sec(\theta)}\)

OpenStudy (anonymous):

@fari321 I did the same.....

OpenStudy (anonymous):

yeah that is just about all you can do

OpenStudy (anonymous):

hmmmm